Circular No. 7970

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 R3 (LONEOS) An apparently asteroidal object reported by LONEOS (discovery ob-

servation below) was posted on the NEO Confirmation Page due to its unusual motion. CCD frames taken on Sept. 7.4 UT by D. T. Durig, Sewanee, TN, with a 0.3-m reflector showed a hint of diffuseness perhaps 8′′–10′′ in size on most of his images; his images on Sept. 5.3 and 12.4 also showed a “soft” appearance, the latter images hinting at a faint tail in p.a. 45^{o}. CCD images taken on Sept. 8.1 at Kleť by M. Tichý and J. Tichá with the 1.06-m KLENOT telescope show the object as a slightly diffuse, faint coma of diameter 12′′, and on Sept. 10.04 it was slightly diffuse with a faint coma of diameter ∼ 12′′–14′′; it appeared diffuse also on images from Sept. 13.03. At Modra, S. Gajdos and A. Galád found a coma diameter of ∼ 5′′ on CCD images taken on Sept. 12.1 (0.6-m reflector). Images taken with the Palomar 1.5-m reflector by M. Hicks on Sept. 14.4 clearly show a 5′′ coma and a slight elongation to the west in a 10-min r-band exposure.

2002 UT  α_{2000}  δ_{2000}  m_{2}
Sept.4.40428  4^{h}25^{m}04.^{s}71  +21^{o}17′58′′.5  18.2

The available observations, parabolic orbital elements (T = 2003 June 14.0 TT, q = 3.864 AU, ω = 45^{o}.2, Ω = 54^{o}.3, i = 161^{o}.1, equinox 2000.0), and an ephemeris appear on MPEC 2002-R75.

NOVA IN M31 K. Hornoch, Lelekovice, Czech Republic, reports his discovery, on R-

band CCD images taken on Aug. 4.013 UT, of an apparent nova (mag 17.0) located at α = 0^{h}42^{m}30^{s}.9, δ = +41^{o}06′13′′ (equinox 2000.0). Nothing is present at this position to mag 20.5 on red Digitized Sky Survey images from 1986 and 1989. Additional R magnitudes from Hornoch (‘K’ denotes photometry from exposures by P. Kušnirák at Ondřejov): 2001 Aug. 16.958, 17.9? (hint or flaw at the limit); 2002 Aug. 17.952, 17.1; 25.859, 18.1; 26.850, 18.0 (K); 30.007, 18.5 (K); 30.867, 18.6; 31.061, 18.6 (K); Sept. 2.951, 18.7. R. J. Foley, S. Jha, M. Papenkova, and A. V. Filippenko, University of California, Berkeley, report that CCD spectra (range 310–1000 nm), obtained on Sept. 14 UT with the Shane 3-m reflector at Lick Observatory, show that this object is indeed a nova, with prominent hydrogen Balmer emission lines and some evidence for weak Fe II emission. Most striking, however, is the unusually strong O I 844.6-nm emission (generally a fluorescence line); its intensity is ∼ 50 percent that of Hα.

2002 September 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green